A sample of a brown-colored gas that is a major air pollutant is found to contain 2.34 g of N and 5.34 g of O. What is the simplest formula of the compound?
1
Expert's answer
2020-06-16T14:58:53-0400

1.We calculate the amount (mol) of 2.34 g of N and the amount (mol) of 5.34 g of O:

n(N)= 2.34/14=0.167 mol

n(O)=5.34/16=0.334 mol

2.We calculate the ratio of O and N:

0.334/0.167=2 so, the formula of the gas is NO2

Answer: NO2